xx CarpetTracker

CarpetTracker

Erik Schnetter <schnetter@cct.lsu.edu>

February 20, 2006

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11surface 10Scope: privatex01 INT

11Description: Spherical surface index which is the source for the location of the refine regions

11 Range 10 x01 Default: -1
11-1  10disabled
110:*  10active

11

11surfacexname 10Scope: privatex01  STRING

11Description: Spherical surface name which is the source for the location of the refine regions

11 Range 10 x01 Default: (none)
11           10use surface
11.*  10any string

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Tell what is going on

11           10  x01  Default: no

11

11nsurfaces 10Scope: shared from SPHERICALSURFACEx01INT

11

7 Interfaces

General Implements: carpettracker

Inherits: sphericalsurface carpetregrid2

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetTracker. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPREREGRID x carpettrackerxsetpositions x set positions of refined regions x

11x11Language:11c
11 x 11Options: 11global
11 x 11Reads: 11sphericalsurface::sfxinfo
11 x 11x  11sphericalsurface::sfxactive
11 x 11x  11sphericalsurface::sfxvalid
11 x 11Type: 11function
11 x 11Writes: 11carpetregrid2::active(everywhere)
11 x 11x  11carpetregrid2::positions(everywhere)